Grade 12Algebra

If a line makes an angle of /4 p with the positive directions of each of x- axis and y- axis, then the angle that the line makes with the positive direction of the z-axis is ?

sudhanshu
12 Years agoGrade 12

1 Answer

SHAIK AASIF AHAMED
12 Years ago
Hello student,
A line makes an angle of \small \pi/4 with thepositivedirections of each of x- axis and y- axis therefore\small \alpha =\pi /4,\small \beta =\pi /4
We know that cos2\small \alpha+cos2\small \beta+cos2\small \gamma=1
substituting the values of\small \alpha and \betain above equation we get
\small \gamma=90
Hence the line makes an angle of 90 with z axis
Thanks and Regards
Shaik Aasif
askIITians faculty